the volume of a pyramid is 108cm3 .the pyramid is 6ft tall.what is the area of the base of the pyramid?please show your work so I can understand.thanks
Answer by sachi(548)

volume of the pyramid with height h & base area a is (1/3)*a*h=ah/3
so ah/3=108 & h=6
so a=108*3/6=54 ft sqrd